# Hierarchical Consensus-Based Multi-Agent Reinforcement Learning for Multi-Robot Cooperation Tasks

### v1

@@ I INTRODUCTION
Multi-Agent Reinforcement Learning (MARL) is garnering increasing attention for its capability to tackle complex tasks [1]. Tasks involving distributed multi-robots [2] often require several agents to collaborate based on their local observations to accomplish a given objective. This requirement aligns with the commonly adopted MARL framework of Centralized Training with Decentralized Execution (CTDE), as exemplified by methods such as MADDPG (Multi-Agent Deep Deterministic Policy Gradient) [3] and MAPPO (Multi-Agent Proximal Policy Optimization) [4]. These approaches utilize global information during the training phase through the critic, while the actor relies solely on individual observations during execution. This setup leads to a significant challenge: agents lack a consensus during task execution, hindering their collective performance for cooperation [5].
In addressing this issue, three primary methodologies have been advanced. The first involves strategies for communication-based multi-agent reinforcement learning [6], which are faced with challenges in selective information sharing and increased bandwidth requirements. The second method leverages intrinsic rewards to create leader-follower dynamics [7], yet this approach is constrained by its task-specific effectiveness and encounters difficulties with broad applicability and generalization. The third approach, employing mean field theory [8], offers a promising direction but often struggles to effectively handle complex tasks. Given existing methods’ limitations, sociology’s research [9, 10] on consensus team agents interacting to align on shared values offers potential solutions for CTDE’s challenges with partial observations.
@@ II-A Consensus in Multi-Agent System
Consensus means the interaction between groups of agents in a team to reach an agreement on a common value or state [13, 14]. Consensus in Multi-Agent Systems has garnered extensive research interest, primarily focusing on achieving shared agreement among agents. Research in this area spans three main categories: first, studies inspired by biological mechanisms and wildlife collective behaviors [15, 16]; second, theoretical explorations using models like the graph theory for foundational insights into consensus [17, 18]; and third, practical applications, including the development of consensus models and protocols [19], with a keen focus on convergence, equilibrium, and implementation challenges. Notably, Multi-Agent Reinforcement Learning (MARL) [4, 20, 21] has become a focal point within the third category, emphasizing the utilization of consensus for enhancing agent cooperation in complex environments.
@@ II-B Contrastive Learning
Recent studies have increasingly leveraged self-supervised learning [22] to enhance model capabilities for downstream tasks. Among these, contrastive learning [23] stands out as a particularly tractable approach. It operates on the principle of minimizing the distance between augmented versions of the same sample while maximizing the distance between distinct ones, as noted by Wang [24]. This method effectively boosts the model’s understanding of equivalent data representations. However, the collection of positive and negative samples in contrastive learning presents challenges that significantly impact its effectiveness. In reinforcement learning (RL) [25], samples are collected through ongoing interactions with the environment, highlighting the potential benefits of integrating contrastive learning with RL. Recent research efforts [26], have successfully employed contrastive learning for representation learning prior to reinforcement learning, achieving unparalleled data efficiency in pixel-based RL tasks. Moreover, contrastive learning has been utilized to formulate reward functions within RL systems [27].
@@ III-A Problem Formulation
The multi-robot cooperation task can be formulated as a decentralized partially observable Markov decision process (Dec-POMDP) [31], defined as $(I,S,A,T,R,O,Z,\gamma)$. The index set $(I=\{1,..,N\})$ represents the set of agents. $S$ is the global state space. Note that each agent is only capable of partial observation of the environment $s\in S$, and the individual observation $o_{i}\in O$ comes from the local observation function $o_{i}=\mathcal{Z}_{i}\left(s\right):S\rightarrow O$. Each agent $i$ chooses its action according to its policy $a_{i}\sim\pi_{i}\left(\cdot\mid o_{i}\right)$. The joint action space $A$ consists of the union of all agent’s action space $\bigcup_{i=1}^{N}\mathcal{A}_{i}$. We define the state transition function $T:S\times A\rightarrow S$ and the discount factor $\gamma\in(0,1)$. All agents share the same joint reward function $R(s,a)$. The agents aim to maximize the expected joint return, defined as $\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R\left(s_{t},a_{t}\right)\right]$
@@ IV-A Consensus Builder
In multi-agent reinforcement learning, agents execute actions based on local observations, leading to a lack of global information guidance during execution. This section focuses on building an effective consensus within the Centralized Training with Decentralized Execution (CTDE) framework. As discussed in a previous section, the DINO framework processes a sample and its data-augmented equivalent through a teacher-student network. We treat an agent’s observation $o$ as the sample, and observations from other agents as the equivalent samples. In other words, the observation function $Z$ is considered an augmentation operation, where the global observation $s$ is augmented into each agent’s observation $o$. These observations $o$ correspond to the same global state $s$.
@@ IV-C HC-MARL Framework
As illustrated in Fig. 4, the hierarchical consensus mechanism enables us to derive the attention-weighted consensus, $c^{att}_{i}$. This consensus serves as the agent’s inferred understanding of the global state, derived from partial observations. We integrate this consensus as an augmented observation input within the multi-agent reinforcement learning framework. It’s pivotal to emphasize that while information from other agents is leveraged during the training phase of the student network, the action execution phase exclusively relies on an agent’s local observations. This design principle ensures that our HC-MARL approach is compatible with a wide range of MARL algorithms, adhering to the Centralized Training with Decentralized Execution (CTDE) paradigm.
@@ V-A Environment Settings
We constructed three cooperative multi-agent tasks within the Webots simulation [33], including Predator-Prey, Rendezvous [34], and Navigation, as shown in Fig. 8. We implemented our HC-MARL in these three tasks and compared it against two main-stream MARL baselines Multi-Agent Proximal Policy Optimization (MAPPO) [4] and its variant Heterogeneous-Agent Proximal Policy Optimization (HAPPO) [35]. Note that our HC-MARL framework can seamlessly integrate with various MARL algorithms. To ensure fair comparisons with these variants of MAPPO, we constructed our HC-MARL framework based on the MAPPO architecture in this experiment.
@@ V-C Ablation Study
In the HC-MARL framework, we employ a hierarchical consensus mechanism to derive short-term and long-term consensus. To assess the efficacy of both the consensus mechanism and the hierarchical approach, we conducted ablation studies varying the number of consensus categories and consensus layers.
@@ V-D Real World Experiments
We validated the real-world applicability of HC-MARL by conducting experiments on E-puck swarm. We utilized the NOKOV motion capture system for indoor positioning, which can be extended to GPS or other visual positioning methods.
@@ VI Conclusion
We introduced the Hierarchical Consensus-Based Multi-Agent Reinforcement Learning (HC-MARL) framework, a novel approach that employs hierarchical consensus to facilitate cooperative execution among agents based on local observations. Recognizing that each agent’s local observations are subsets of a consistent global state, our framework leverages contrastive learning from these observations to achieve a global consensus, which then serves as additional local observations for the agents. By implementing a hierarchical mechanism, we construct short-term and long-term consensus to cater to the dynamic requirements of various tasks. This process is further refined through an adaptive attention mechanism that dynamically adjusts the influence of each consensus layer. Extensive experiments demonstrate that the HC-MARL method significantly enhances the performance of multi-robot cooperation tasks.
@@ VII ACKNOWLEDGMENT
We gratefully acknowledge the support of the National Key R&D Program of China (2022ZD0116401) and the National Natural Science Foundation of China (Grant No. 62306023).

### v2

@@ I INTRODUCTION
Multi-Agent Reinforcement Learning (MARL) is garnering increasing attention for its capability to tackle complex tasks [1, 2]. Tasks involving distributed multi-robots [3, 4] often require several agents to collaborate based on their local observations to accomplish a given objective. This requirement aligns with the commonly adopted MARL framework of Centralized Training with Decentralized Execution (CTDE), as exemplified by methods such as MADDPG (Multi-Agent Deep Deterministic Policy Gradient) [5] and MAPPO (Multi-Agent Proximal Policy Optimization) [6]. These approaches utilize global information during the training phase through the critic, while the actor relies solely on individual observations during execution. This setup presents a substantial challenge: agents fail to reach a consensus during task execution, which impedes their collective efficacy in cooperation.
In addressing this issue, three primary methodologies have been advanced. The first involves strategies for communication-based multi-agent reinforcement learning [7], which are faced with challenges in selective information sharing and increased bandwidth requirements. The second method leverages intrinsic rewards to create leader-follower dynamics [8], yet this approach is constrained by its task-specific effectiveness and encounters difficulties with broad applicability and generalization. The third approach, employing mean field theory [9], offers a promising direction but often struggles to effectively handle complex tasks. Given existing methods’ limitations, sociology’s research [10] on consensus team agents interacting to align on shared values offers potential solutions for CTDE’s challenges with partial observations.
@@ II-A Consensus in Multi-Agent System
Consensus means the interaction between groups of agents in a team to reach an agreement on a common value or state [13]. Consensus in Multi-Agent Systems has garnered extensive research interest, primarily focusing on achieving shared agreement among agents. Research in this area spans three main categories: first, studies inspired by biological mechanisms and wildlife collective behaviors [14, 15]; second, theoretical explorations using models like the graph theory for foundational insights into consensus [16, 17]; and third, practical applications, including the development of consensus models and protocols [18], with a keen focus on convergence, equilibrium, and implementation challenges. Notably, Multi-Agent Reinforcement Learning (MARL) [19, 20, 21] has become a focal point within the third category, emphasizing the utilization of consensus for enhancing agent cooperation in complex environments.
@@ II-B Contrastive Learning
Recent studies have increasingly leveraged self-supervised learning [22] to enhance model capabilities for downstream tasks. Among these, contrastive learning [23] stands out as a particularly tractable approach. It operates on the principle of minimizing the distance between augmented versions of the same sample while maximizing the distance between distinct ones, as noted by Wang [24]. This method effectively boosts the model’s understanding of equivalent data representations. However, the collection of positive and negative samples in contrastive learning presents challenges that significantly impact its effectiveness. In reinforcement learning (RL), samples are collected through ongoing interactions with the environment, highlighting the potential benefits of integrating contrastive learning with RL. Recent research efforts [25], have successfully employed contrastive learning for representation learning prior to reinforcement learning, achieving unparalleled data efficiency in pixel-based RL tasks. Moreover, contrastive learning has been utilized to formulate reward functions within RL systems [26].
@@ III-A Problem Formulation
The multi-robot cooperation task can be formulated as a decentralized partially observable Markov decision process (Dec-POMDP), defined as $(I,S,A,T,R,O,Z,\gamma)$. The index set $(I=\{1,..,N\})$ represents the set of agents. $S$ is the global state space. Note that each agent is only capable of partial observation of the environment $s\in S$, and the individual observation $o_{i}\in O$ comes from the local observation function $o_{i}=\mathcal{Z}_{i}\left(s\right):S\rightarrow O$. Each agent $i$ chooses its action according to its policy $a_{i}\sim\pi_{i}\left(\cdot\mid o_{i}\right)$. The joint action space $A$ consists of the union of all agent’s action space $\bigcup_{i=1}^{N}\mathcal{A}_{i}$. We define the state transition function $T:S\times A\rightarrow S$ and the discount factor $\gamma\in(0,1)$. All agents share the same joint reward function $R(s,a)$. The agents aim to maximize the expected joint return, defined as $\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R\left(s_{t},a_{t}\right)\right]$
@@ IV-A Consensus Builder
In multi-agent reinforcement learning, agents execute actions based on local observations, leading to a lack of global information guidance during execution. Inspired by Xu [28], we utilized insights from his method to develop a consensus mechanism based on local observations. This section focuses on building an effective consensus within the CTDE framework. As discussed in a previous section, the DINO framework processes a sample and its data-augmented equivalent through a teacher-student network. We treat an agent’s observation $o$ as the sample, and observations from other agents as the equivalent samples. In other words, the observation function $Z$ is considered an augmentation operation, where the global observation $s$ is augmented into each agent’s observation $o$. These observations $o$ correspond to the same global state $s$.
@@ IV-C HC-MARL Framework
As illustrated in Fig. 4, the hierarchical consensus mechanism enables us to derive the attention-weighted consensus, $c^{att}_{i}$. This consensus serves as the agent’s inferred understanding of the global state, derived from partial observations. We integrate this consensus as an augmented observation input within the multi-agent reinforcement learning framework. It’s pivotal to emphasize that while information from other agents is leveraged during the training phase of the student network, the action execution phase exclusively relies on an agent’s local observations. This design principle ensures that our HC-MARL approach is compatible with a wide range of MARL algorithms, adhering to the CTDE paradigm.
@@ V-A Environment Settings
We constructed three cooperative multi-agent tasks within the Webots simulation, including Predator-Prey, Rendezvous, and Navigation, as shown in Fig. 8. We implemented our HC-MARL in these three tasks and compared it against two main-stream MARL baselines Multi-Agent Proximal Policy Optimization (MAPPO) and its variant Heterogeneous-Agent Proximal Policy Optimization (HAPPO) [31]. Note that our HC-MARL framework can seamlessly integrate with various MARL algorithms. To ensure fair comparisons with these variants of MAPPO, we constructed our HC-MARL framework based on the MAPPO architecture in this experiment.
@@ V-C Ablation Study
To assess the efficacy of both the consensus mechanism and the hierarchical approach, we conducted ablation studies varying the number of consensus categories and layers.
@@ V-D Real World Experiments
We validated the real-world applicability of HC-MARL by conducting experiments on E-puck swarm. We utilized the NOKOV motion capture system for indoor positioning.
@@ VI Conclusion
We introduced the Hierarchical Consensus-Based Multi-Agent Reinforcement Learning (HC-MARL) framework, a novel approach that employs hierarchical consensus to facilitate cooperative execution among agents based on local observations. Recognizing that each agent’s local observations are subsets of a consistent global state, our framework leverages contrastive learning from these observations to achieve a global consensus, which then serves as additional local observations for the agents. By implementing a hierarchical mechanism, we construct short-term and long-term consensus to cater to the dynamic requirements of various tasks. Extensive experiments demonstrate that the HC-MARL method significantly enhances the performance of multi-robot cooperation tasks.
@@ VII ACKNOWLEDGMENT
We gratefully acknowledge the support of the National Science and Technology Major Project(No. 2022ZD0116401), the National Natural Science Foundation of China (Grant No. 62306023), and the Science and Technology Project of State Grid Corporation of China (No. 5108-202218280A-2-402-XG).